# DC-Gaussian: Improving 3D Gaussian Splatting for Reflective Dash Cam Videos

### v2

@@ Abstract
We present DC-Gaussian, a new method for generating novel views from in-vehicle dash cam videos. While neural rendering techniques have made significant strides in driving scenarios, existing methods are primarily designed for videos collected by autonomous vehicles. However, these videos are limited in both quantity and diversity compared to dash cam videos, which are more widely used across various types of vehicles and capture a broader range of scenarios. Dash cam videos often suffer from severe obstructions such as reflections and occlusions on the windshields, which significantly impede the application of neural rendering techniques. To address this challenge, we develop DC-Gaussian based on the recent real-time neural rendering technique 3D Gaussian Splatting (3DGS). Our approach includes an adaptive image decomposition module to model reflections and occlusions in a unified manner. Additionally, we introduce illumination-aware obstruction modeling to manage reflections and occlusions under varying lighting conditions. Lastly, we employ a geometry-guided Gaussian enhancement strategy to improve rendering details by incorporating additional geometry priors. Experiments on self-captured and public dash cam videos show that our method not only achieves state-of-the-art performance in novel view synthesis, but also accurately reconstructing captured scenes getting rid of obstructions. Project page: https://linhanwang.github.io/dcgaussian/.
@@ 2.1 Novel view synthesis for driving scenes
Recently 3DGS [19] has attracted great attention in the research community. It achieves optimal results in novel view synthesis and rendering speed by explicitly modeling a 3D scene with 3D gassians. Some researchers have extended it to dynamic objects and scenes. Given a set of dynamic monocular images, a deformation network is introduced to model the motion of Gaussians [56]. DrivingGaussian [65] models and decomposes dynamic driving scenes based on composite gaussian splatting. GaussianPro [9] improves the geometry of 3DGS by controlling the densification process of 3DGS with classic PatchMatching algorithm [2]. Zhou et al. [64] proposed to utilize 3D Gaussian Splatting for holistic urban scene understanding. However, dash cam videos, an important data source for understanding driving scenes, remain unexplored due to obstructions on the windshield. Despite previous works [65, 52, 56] making progress in separating dynamic objects from background scenes, the image decomposition problem we are addressing presents unique challenges because of the obstructions’ transparent or semi-transparent nature.
@@ 2.2 Obstruction removal and layer separation
Multi-image layer separation. Existing methods often exploit differences in motion patterns between transmission and obstruction layers and use learned image priors [15] to decompose images into multiple components. These layer separation methods estimate dense motion fields for each layer using optical flow [42], SIFT flow [39], and deep learning-based flow estimation methods [40, 27]. Recently, Nam et al. [31] proposed a multi-frame fusion framework based on neural image representation, achieving strong performance on various layer-separation tasks. Similarly, NSF [11] fuses RAW burst image sequences by modeling optical flow with neural spline fields. However, methods designed for burst images struggle with large pixel motions in driving scenes.
@@ 3.2 Adaptive image decomposition
$I(u,v,j)=(1-\boldsymbol{\phi})I_{t}(u,v,j)+I_{o}(u,v,j)$ (4)
where $u,v\in[0,1]$ are the continuous image coordinates and $j\in[0,1,...,N-1]$ is the frame number. Instead of defining the opacity field in 3D world space, we define the opacity relative to the 2D image space of each view. This approach is more convenient for modeling obstructions because view-dependent effects are only related to the training images. In this image decomposition method, the transmission images represent the driving scenes viewed through the windshield. We can use standard 3DGS to model these scenes due to its multi-view consistent property. Thus, $I_{t}(u,v,j)$ can be easily calculated by Eq. 2, given the camera pose of frame $j$. $I_{o}$ represents the appearance of the complex obstructions on the windshield. We explain its modeling in the next section.
@@ 3.3 Illumination-aware Obstruction Modeling
Figure 4: When the intensity of incident light changes, the strength of reflections also change accordingly (a)(d). Our method achieves high-fidelity reflections synthesis (c)(f) and reasonable decomposition results (b)(e) under varying light. The reflections in (f) are brightened to reveal details.
@@ 3.4 Geometry-guided Gaussian Enhancement
In IOM, we utilize the motion pattern prior of obstructions to reduce ambiguity in decomposing images. Despite this, some ambiguity still remains. To further enhance performance, we incorporate geometry priors. Typically, strong obstructions affect only portions of the images. As cars move, the same areas in the 3D world can appear in several image fragments which are not or less interfered with by obstructions. In these image fragments, capturing the same areas from different angles, ambiguity is minimal due to the dominance of transmission. Based on this intuition, we leverage a multi-view stereo (MVS) algorithm to identify these "image fragments" and generate geometry priors for 3D Gaussians.
Specifically, we employ a deep MVS method [47] to generate depth maps for all views. A geometric consistency filtering process [36] is leveraged to generate masks $M_{j}$ masking the multi-view consistent areas, which are essentially the "image fragments" we want to identify. The masked depth maps are then mapped to 3D space as dense point clouds, which are used to initialize 3D Gaussians at physically accurate positions. To find more multi-view consistent "image fragments," we propose suppressing the obstructions in training images by employing our proposed method. Specifically, we remove obstructions from input images to obtain transmission $\hat{I}^{j}_{t}$. For areas blocked by occlusions, we inpaint the content with $I^{j}_{t}$.
@@
Here $I^{j}$ is the $j_{th}$ input image. $I^{j}_{o}$ and $I^{j}_{t}$ are synthesized by trained IOM and 3DGS, respectively. We use 0.5 for the threshold $\tau$ in all the experiments.
@@ 3.5 Implementation details
We develop our method based on 3DGS. We borrow multi-resolution hash encoding and fast MLP implementation from tiny-cuda-nn [30] to build IOM. We choose PatchMatchNet [47] as the MVS method in G3E. We follow previous driving scenes reconstruction works [45, 10] to separately model the sky areas. We use a $L_{1}$ loss $\mathcal{L}_{opacity}$ to regularize the opacity field . The final loss function is:
@@
$\mathcal{L}_{pho}$ is the same as in standard 3DGS [19]. $\mathcal{L}_{sky}$ is borrowed from UCNeRF [10]. We use $0.001$ for both $\lambda_{1}$ and $\lambda_{2}$. We run all the experiments with an A100 GPU. Each scene contains approximately 300 images in our evaluation datasets. Combined training of the 3DGS and IOM for 30k iterations with Adam optimizer takes about 30 minutes. It takes 40 minutes in total to evaluate MVS and run geometry filtering.
Figure 5: Comparisons with 3DGS on novel view synthesis. Because the obstructions violate multi-view consistency, the performance of 3DGS degrades significantly, resulting in artifacts and blurry renderings (highlighted by red arrows). In contrast, our method not only faithfully synthesizes novel view renderings but also renders transmission with fine details.
@@ 4.1 Datasets
BDD100K. To evaluate the performance of our method and baselines, we adopt BDD100K [57] for evaluation. These scenes are from dash cam videos captured in daily life. They contain common obstructions, such as reflections, mobile phone holders, stickers and stains. Evaluation on this dataset reflects performance on real life dash cam videos.
DCVR. To further evaluate the performance of our method on strong reflection conditions, we established DCVR (Dash Cam Videos with Reflection) dataset. The original videos are undistorted [5] to ease the structure-from-motion algorithm. We utilize the popular tools COLMAP [36] and HLoc [34, 26] to estimate the camera parameters.
@@ 4.2 Baselines
We choose 3DGS as our baseline [19]. We also compare our method with other state-of-the-art method Zip-NeRF [4] and GaussianPro [9]. To evaluate the performance of novel view synthesis, following common settings [3], we select one of every eight images as testing images and the remaining ones for training. Since our method is also designed for image decomposition, we also compare our method with state-of-the-art obstruction removal methods, including DSRNet [17], NIR [31] and Liu et al. [27].
@@ 4.3 Quantitative results
For the quantitative evaluation, we conduct comparison with baselines on both BDD100K and DCVR. We apply the three widely-used metrics for evaluation, i.e., PSNR, SSIM [48], and LPIPS [60]. The results are shown in Tab. 1. The consistent superior performance of our method shows the efficacy of the proposed modules. Though GaussianPro applies a progressive propagation strategy to enhance the geometry of 3D Gasssians, without separate obstruction modeling, it cannot handle the obstructions corrupted dash cam videos.
               (a) Reference                               (b) Transmission                             (c) Obstruction
                 (d) NIR [31]                                (e) DSRNet [17]                           (f) Liu et al. [27]
Figure 6: Comparisons with the image-based reflection removal methods on removing reflections. (a) Reference is a frame in a dash cam video. (d), (e), and (f) are results from previous obstruction removal methods, which are not effective in this scenario. In comparison, our method decomposes the image and synthesizes (b) transmission and (c) obstruction with high fidelity.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
ZipNeRF [4] 27.89 0.875 0.176 24.41 0.786 0.228
GaussianPro [9] 27.75 0.894 0.192 23.71 0.770 0.270
3DGS [19] 28.02 0.897 0.188 23.73 0.783 0.248
DCGaussian (Ours) 29.44 0.914 0.143 24.74 0.822 0.202
@@
Figure 7: Ablation study about the Learnable opacity map $\boldsymbol{\phi}$. The incorporation of the opacity map leads to better view synthesis and obstruction removal.
@@ 5 Conclusions
In conclusion, we propose DC-Gaussian that effectively addresses the challenges of extending 3D Gaussian Splatting to dash cam videos. Experiments on BDD100K and DCVR demonstrate a significant improvement in rendering quality and images decomposition, setting a new benchmark for neural rendering with dash cam videos.
Limitations and future work Currently, DC-Gaussian has only been evaluated on single-sequence videos. However, considering the vast amount of dash cam footage available, extending DC-Gaussian to a multi-sequence video setting and leveraging dense view images to achieve more pleasing results would be a promising direction for future research.
@@ Appendix A Appendix / supplemental material
(a) Reference
(b) Transmission
(c) Obstruction
(d) DSRNet [17]
(e) NIR [31]
(f) Liu et al. [27]
Figure 11: Comparisons with image-based reflection removal methods show that our method achieves significantly better results than previous methods [17, 31, 27] in removing reflections.

### v3

@@ Abstract
We present DC-Gaussian, a new method for generating novel views from in-vehicle dash cam videos. While neural rendering techniques have made significant strides in driving scenarios, existing methods are primarily designed for videos collected by autonomous vehicles. However, these videos are limited in both quantity and diversity compared to dash cam videos, which are more widely used across various types of vehicles and capture a broader range of scenarios. Dash cam videos often suffer from severe obstructions such as reflections and occlusions on the windshields, which significantly impede the application of neural rendering techniques. To address this challenge, we develop DC-Gaussian based on the recent real-time neural rendering technique 3D Gaussian Splatting (3DGS). Our approach includes an adaptive image decomposition module to model reflections and occlusions in a unified manner. Additionally, we introduce illumination-aware obstruction modeling to manage reflections and occlusions under varying lighting conditions. Lastly, we employ a geometry-guided Gaussian enhancement strategy to improve rendering details by incorporating additional geometry priors. Experiments on self-captured and public dash cam videos show that our method not only achieves state-of-the-art performance in novel view synthesis, but also accurately reconstructing captured scenes getting rid of obstructions. See the project page for code, data: https://linhanwang.github.io/dcgaussian/.
@@ 2.1 Novel View Synthesis for Driving Scenes
Recently 3DGS [19] has attracted great attention in the research community. It achieves optimal results in novel view synthesis and rendering speed by explicitly modeling a 3D scene with 3D gaussians. Some researchers have extended it to dynamic objects and scenes. Given a set of dynamic monocular images, a deformation network is introduced to model the motion of Gaussians [58]. DrivingGaussian [67] models and decomposes dynamic driving scenes based on composite gaussian splatting. GaussianPro [9] improves the geometry of 3DGS by controlling the densification process of 3DGS with classic PatchMatching algorithm [2]. Zhou et al. [66] propose to utilize 3D Gaussian Splatting for holistic urban scene understanding. However, dash cam videos, an important data source for understanding driving scenes, remain unexplored due to obstructions on the windshield. Despite previous works [67, 54, 58] making progress in separating dynamic objects from background scenes, the image decomposition problem we are addressing presents unique challenges because of the obstructions’ transparent or semi-transparent nature.
@@ 2.2 Obstruction Removal and Layer Separation
Multi-image layer separation. Existing methods often exploit differences in motion patterns between transmission and obstruction layers and use learned image priors [15] to decompose images into multiple components. These layer separation methods estimate dense motion fields for each layer using optical flow [43], SIFT flow [40], and deep learning-based flow estimation methods [41, 27]. Recently, Nam et al. [32] propose a multi-frame fusion framework based on neural image representation, achieving strong performance on various layer-separation tasks. Similarly, NSF [11] fuses RAW burst image sequences by modeling optical flow with neural spline fields. However, methods designed for burst images struggle with large pixel motions in driving scenes.
@@ 3.2 Adaptive Image Decomposition
$I(u,v,j)=(1-\boldsymbol{\phi})I_{t}(u,v,j)+\boldsymbol{\phi}*I_{o}(u,v,j)$ (4)
where $u,v\in[0,1]$ are the continuous image coordinates and $j\in[0,1,...,N-1]$ is the frame index. Instead of defining the opacity field in 3D world space, we define the opacity relative to the 2D image space of each view, resulting in a learnable 2D tensor in practice. This approach is more convenient for modeling obstructions because view-dependent effects are only related to the training images. In this image decomposition method, the transmission images represent the driving scenes viewed through the windshield. We can use standard 3DGS to model these scenes due to its multi-view consistent property. Thus, $I_{t}(u,v,j)$ can be easily calculated by Eq. 2, given the camera pose of frame $j$. $I_{o}(u,v,j)$ represents the appearance of the complex obstructions on the windshield. In the implementation, we incorporate $\phi$ into $I_{o}(u,v,j)$ in the second term of Eq. 4 for robust training. We explain the modeling of $I_{o}(u,v,j)$ in the next section.
@@ 3.3 Illumination-aware Obstruction Modeling
Figure 4: When the intensity of incident light changes, the strength of reflections also changes accordingly (a, d). Our method achieves high-fidelity reflections synthesis (c, f) and reasonable decomposition results (b, e) under varying light. The reflections in (f) are too weak to be seen by the eye, so we brighten it to reveal the details.
@@ 3.4 Geometry-Guided Gaussian Enhancement
In IOM, we utilize the motion pattern prior of obstructions to reduce ambiguity in decomposing images. Despite this, some ambiguity still remains. To further enhance performance, we incorporate geometry priors. Typically, strong obstructions affect only portions of the images. As cars move, the same objects in the 3D world can appear in several image fragments which are not or less interfered with by obstructions. In these image fragments, the texture in the transmission is less blurred by obstructions, and multi-view consistency is maintained. Based on this intuition, we leverage a multi-view stereo (MVS) algorithm to identify these "image fragments" and generate geometry priors for 3D Gaussians.
Specifically, we employ a deep MVS method [48] to generate depth maps for all views. A geometric consistency filtering process [37] is leveraged to generate masks $M_{j}$ masking the multi-view consistent areas, which are essentially the "image fragments" we want to identify. The masked depth maps are then mapped to 3D space as dense point clouds, which are used to initialize 3D Gaussians at physically accurate positions. To find more multi-view consistent "image fragments," we propose suppressing the obstructions in training images by employing our proposed method. Specifically, we remove obstructions from input images to obtain transmission $\hat{I}^{j}_{t}$. For areas blocked by occlusions(where $\boldsymbol{\phi}$ has large value), we inpaint the content with $I^{j}_{t}$.
@@
Here $I^{j}$ is the $j_{th}$ input image. $I^{j}_{o}$ and $I^{j}_{t}$ are synthesized by trained IOM and 3DGS, respectively. Eq. 7 is derived from Eq. 4. We use 0.5 for the threshold $\tau$ in all the experiments.
@@ 3.5 Implementation Details
We develop our method based on 3DGS. We borrow multi-resolution hash encoding and fast MLP implementation from tiny-cuda-nn [31] to build IOM. We choose PatchMatchNet [48] as the MVS method in G3E. We follow previous driving scenes reconstruction works [46, 10] to separately model the sky areas. The final loss function is:
@@
$\mathcal{L}_{pho}$ is the same as in standard 3DGS [19]. $\mathcal{L}_{sky}$ is borrowed from UCNeRF [10]. We use a $L_{1}$ loss $\mathcal{L}_{opacity}=\sum_{(u,v)}\|\phi(u,v)\|_{1}$ to regularize the opacity field, where $(u,v)$ are the image coordinates. This opacity loss encourages the opacity map to have the minimum areas that could satisfy the optimization. This design is based on the prior knowledge that opaque objects typically occupy only small portions of the windshield. We use $0.001$ for both $\lambda_{1}$ and $\lambda_{2}$. We run all the experiments with an A100 GPU. Each scene contains approximately 300 images in our evaluation datasets. Combined training of the 3DGS and IOM for 30k iterations with Adam optimizer takes about 30 minutes. It takes 40 minutes in total to evaluate MVS and run geometry filtering.
Figure 5: Comparisons with 3DGS on novel view synthesis. Because the obstructions violate multi-view consistency, the performance of 3DGS degrades significantly, resulting in artifacts and blurry renderings (highlighted by red arrows). In contrast, our method not only faithfully synthesizes novel view renderings but also renders transmission with fine details, exhibiting an improvement of 3.05dB in terms of PSNR.
@@ 4.1 Datasets
BDD100K. To evaluate the performance of our method and baselines, we adopt BDD100K [59] for evaluation. This dataset contains 8 scenes that are from dash cam videos captured in daily life. They contain common obstructions, such as reflections, mobile phone holders, stickers and stains. Evaluation on this dataset reflects performance on real life dash cam videos.
DCVR. To further evaluate the performance of our method on strong reflection conditions, we established DCVR (Dash Cam Videos with Reflection) dataset. This dataset contains 10 dash cam videos we collect. The original videos are undistorted [5] to ease the structure-from-motion algorithm. We utilize the popular tools COLMAP [37] and HLoc [35, 26] to estimate the camera parameters.
In both datasets, each sequence consists of approximately 300 frames, extracted from 10-second videos at a frame rate of 30 Hz. For scenes containing car hoods, the lower parts of the images are removed during preprocessing. Seven out of eighteen scenes in two datasets involve car turns, introducing diverse illumination changes. Additional details and visual results are provided in the appendix.
@@ 4.2 Baselines
We choose 3DGS as our baseline [19]. We also compare our method with other state-of-the-art methods Zip-NeRF [4], NeRF-W [29] and GaussianPro [9]. We use the unofficial implementations of Zip-NeRF  [^1] https://github.com/SuLvXiangXin/zipnerf-pytorch and NeRF-W  [^2] https://github.com/kwea123/nerf_pl. To evaluate the performance of novel view synthesis, following common settings [3], we select one of every eight images as testing images and the remaining ones for training. Since our method is also designed for image decomposition, we also compare our method with state-of-the-art obstruction removal methods, including DSRNet [17], NIR [32] and Liu et al. [27].
@@ 4.3 Quantitative Results
For the quantitative evaluation, we conduct comparison with baselines on both BDD100K and DCVR. We apply the three widely-used metrics for evaluation, i.e., PSNR, SSIM [49], and LPIPS [62]. The results are shown in Tab. 1. The consistent superior performance of our method shows the efficacy of the proposed modules. Though GaussianPro and Zip-NeRF achieve great performance in obstruction-free scenes with their progressive propagation strategy and anti-aliasing mechanism, without separate obstruction modeling, they cannot handle the obstructions corrupted dash cam videos. NeRF-W is designed to handle illumination and content difference between images taken at different times but still cannot handle obstructions on the windshield. We show more visual results in the appendix. DC-Gaussian achieves 120 fps at a resolution of 1920x1080 on an RTX 3090 GPU. Although ours is slightly slower than 3DGS, the speed still enables real-time rendering, which is crucial for applications such as autonomous driving simulators.
Figure 6: Comparisons with the image-based reflection removal methods on removing reflections. (a) Reference is a frame in a dash cam video. (b) and (c) are transmission and obstruction decomposed by our method. (d), (e), and (f) are results from previous obstruction removal methods, NIR [32], DSRNet [17] and Liu et al. [27], which are not effective in this scenario. In comparison, our method decomposes the image and synthesizes (b) transmission and (c) obstruction with high fidelity.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
NeRF-W [29] 22.58 0.708 0.395 - - - 0.18
ZipNeRF [4] 27.89 0.875 0.176 24.41 0.786 0.228 0.27
GaussianPro [9] 27.75 0.894 0.192 23.71 0.770 0.270 210
3DGS [19] 28.02 0.897 0.188 23.73 0.783 0.248 155
DCGaussian (Ours) 29.44 0.914 0.143 24.74 0.822 0.202 120
@@ 4.5 Ablation Study
Figure 7: Ablation study about the Learnable opacity map $\boldsymbol{\phi}$. Incorporating the opacity map allows our method to accurately identify the positions of opaque objects, enhancing physical simulation and improving view synthesis and obstruction removal. Without the opacity map, severe artifacts appear.
@@ 5 Conclusions
In conclusion, we propose DC-Gaussian, which effectively addresses the challenges of extending 3D Gaussian Splatting to dash cam videos for the first time. The proposed Adaptive Image Decomposition module enables unified modeling of reflections and occlusions. To handle the reflections and occlusions under challenging lighting conditions, we introduce Illumination-aware Obstruction modeling. Additionally, we employ a Geometry-Guided Gaussian Enhancement strategy to further improve rendering quality. Experiments on BDD100K and DCVR demonstrate significant improvements in rendering quality and image decomposition, setting a new benchmark for neural rendering with dash cam videos.
Limitations and future work Currently, DC-Gaussian has only been evaluated on single-sequence videos. However, considering the vast amount of dash cam footage available, extending DC-Gaussian to a multi-sequence video setting and leveraging dense view images to achieve more pleasing results would be a promising direction for future research. In addition, Our method is not specifically designed to improve performance on dynamic scenes. We provide additional experimental results in the appendix. The results demonstrate that dynamic objects do not significantly impact the performance of obstruction removal. When dynamic objects move at a slow speed, our method also presents reasonable results. We plan to incorporate techniques [53] for dynamic objects modeling into our method in future research to enable robust dynamic modeling.
Acknowledgments and Disclosure of Funding
The authors acknowledge Advanced Research Computing at Virginia Tech for providing computational resources and technical support that have contributed to the results reported within this paper. URL: https://arc.vt.edu/